Doug hits a hockey puck, giving it an initial velcoity of 6.0 m/s. If the coefficient of kinetic friction between ice and puck is 0.050, how far will the puck slide before stopping?
Answer #1

acceleration = F/m = (coef. friction) g

acceleration = 0.050 * 9.81 m/s2


= 0.4905 m/s2

-----------------------------

v = at

6 = 0.4905 t

t = 12.23 s

----------------------------

s = vt - 0.5at^2


= (6 * 12.23) - (0.5 * 0.4905 * 12.23^2)

s = 36.7 m